# Lie algebras generated by reflections in types BCD

### v1

@@ 1.1. The Lie algebra of transpositions
Motivated by questions from the representation theory of the braid group, Marin [13] solved the preceding problem in the case $W=\mathfrak{S}_{n}$ of the symmetric group on $n$ letters (i.e., the Weyl group of type $A_{n-1}$) by computing the structure of the Lie algebra $\mathfrak{s}_{n}$ generated by the set of transpositions in $\mathfrak{S}_{n}$. Recall that the isomorphism classes of the simple $\mathbb{C}\mathfrak{S}_{n}$-modules are labeled by integer partitions of $n$. Marin showed that $\mathfrak{s}_{n}$ is a reductive Lie algebra whose semisimple derived subalgebra $\mathfrak{s}_{n}^{\prime}=[\mathfrak{s}_{n},\mathfrak{s}_{n}]$ is a direct sum of special linear Lie algebras, indexed by equivalence classes of partitions whose Young diagrams are non-symmetric and not hooks, and of orthogonal and symplectic Lie algebras, indexed by partitions whose Young diagrams are symmetric and not hooks. A more detailed summary of Marin’s result is given below in Section 4.3.4. The calculation of $\mathfrak{s}_{n}$ enabled Marin to determine the algebraic hull of the image of the braid group in each of the irreducible representations of the Iwahori–Hecke algebra of type A, and in the Iwahori–Hecke algebra itself. Earlier work by Marin [12] also showed that the computation of $\mathfrak{s}_{n}$ permits one to describe how tensor products of representations for the Iwahori–Hecke algebra decompose as representations over the braid group; see [2, 14] for related and additional results.
@@ 1.2. The Lie algebra of reflections in types BCD
The proof of Theorem 5.1.1 is by induction on $n$, with the values $n\leq 5$ handled via calculations in GAP [6], and the inductive step for $n\geq 6$ treated in Sections 5.25.10. The overall inductive strategy parallels that employed by Marin for type A, but the statements of results and the arguments required in types B and D can be substantially more involved. In particular, types B and D require an intricate, case-by-case analysis of the implications of the branching rules when one restricts simple modules from $\mathbb{C}\mathcal{B}_{n}$ to $\mathbb{C}\mathcal{B}_{n-1}$ and from $\mathbb{C}\mathcal{D}_{n}$ to $\mathbb{C}\mathcal{D}_{n-1}$.
@@ 1.3. Further questions
The results of this paper open the door to investigating analogues for types BCD of the applications pursued by Marin in type A. For example, one can ask for the algebraic hull of the Artin groups of types B and D inside the corresponding characteristic zero Iwahori–Hecke algebras. The image of these Artin groups inside the finite Iwahori–Hecke algebra (defined over a finite field) has been determined by Esterle [4] in the semisimple case.
@@
If $W$ is a finite Coxeter group, then there exists a canonical supergroup structure on $W$ (i.e., a multiplicative grading by $\mathbb{Z}/2\mathbb{Z}$) defined by declaring the Coxeter generators to each be of odd superdegree. This grading extends to the group algebra $\mathbb{C}W$, making it an associative superalgebra. One can then consider the induced Lie superalgebra structure on $\mathbb{C}W$, with Lie bracket given by the graded commutator of elements, and ask for the structure of the Lie superalgebra generated by the set of reflections. In [3] the authors considered this problem in the case $W=\mathfrak{S}_{n}$, and showed that the Lie superalgebra generated by the transpositions is equal to the entire derived Lie superalgebra of $\mathbb{C}\mathfrak{S}_{n}$ plus the span of an odd central element. Generalizing that work to types BCD (and beyond) would be interesting and nontrivial.
@@ 2. Type B classical representation theory
In Sections 2 and 3 we recall some classical results for Weyl groups of types $B_{n}$ and $D_{n}$, many of which can be found in Chapters 1 and 5 of [8].
@@ 2.1. Type B Weyl groups
Let $\mathcal{B}_{n}\subseteq GL_{n}(\mathbb{C})$ be the group of signed permutation matrices, i.e., matrices that have precisely one nonzero entry in each row and each column, equal to either $+1$ or $-1$. Let $H_{n}\subseteq\mathcal{B}_{n}$ be the subgroup of permutation matrices in $\mathcal{B}_{n}$ (i.e., matrices in $\mathcal{B}_{n}$ whose nonzero entries are all $+1$), and let $N_{n}\subseteq\mathcal{B}_{n}$ be the subgroup of diagonal matrices in $\mathcal{B}_{n}$. Then $H_{n}$ identifies with the symmetric group $\mathfrak{S}_{n}$, $N_{n}$ identifies with the elementary abelian $2$-group $\mathbb{Z}_{2}^{n}$, and $\mathcal{B}_{n}=N_{n}\rtimes H_{n}$. Henceforward we may make the identification $\mathcal{B}_{n}=N_{n}\rtimes\mathfrak{S}_{n}$.
For $11\leq i\leq n-$ 1, let $s_{i}\in H_{n}\subseteq\mathcal{B}_{n}$ be the matrix obtained from the identity matrix by interchanging the $i$-th and $(i+1)$-th rows, so $s_{i}$ corresponds to the transposition $(i,i+1)\in\mathfrak{S}_{n}$. For $1\leq i\leq n$, let $t_{i}\in N_{n}\subseteq\mathcal{B}_{n}$ be the diagonal matrix whose $i$-th diagonal entry is equal to $-1$ and whose other diagonal entries are equal to $+1$, so $t_{i}$ corresponds to a multiplicative generator for the $i$-th factor in $\mathbb{Z}_{2}^{n}$. Then for $n\geq 2$, the pair $(\mathcal{B}_{n},\left\{s_{1},\ldots,s_{n-1},t_{n}\right\})$ is a Coxeter system of type $B_{n}$. Each element of $\mathcal{B}_{n}$ is uniquely of the form $t_{1}^{i_{1}}\cdots t_{n}^{i_{n}}\sigma$ with $\sigma\in H_{n}$ and $i_{j}\in\left\{0,1\right\}$ for each $j$. Identifying $H_{n}$ with $\mathfrak{S}_{n}$, the set of reflections in $\mathcal{B}_{n}$ is
@@
The two sets in this union are each conjugacy classes in $\mathcal{B}_{n}$. Following [16, §2] or [15, §2], we may label these conjugacy classes as $C_{(21^{n-2},\emptyset)}$ and $C_{(1^{n-1},1)}$, respectively.
@@ 2.2. Simple modules for Type B Weyl groups
The map $\varepsilon$ is the sign character of $\mathcal{B}_{n}$, i.e., the character that evaluates to $-1$ on each of the Coxeter generators of $\mathcal{B}_{n}$. The map $\varepsilon^{\prime\prime}$ factors via the canonical quotient map $\mathcal{B}_{n}=N_{n}\rtimes H_{n}\twoheadrightarrow H_{n}\cong\mathfrak{S}_{n}$ to the usual the sign character on $\mathfrak{S}_{n}$. By abuse of notation, we identify $\varepsilon$, $\varepsilon^{\prime}$, and $\varepsilon^{\prime\prime}$ with their corresponding one-dimensional $\mathbb{C}\mathcal{B}_{n}$-modules. Then $\alpha=\beta\otimes\gamma$ whenever $\{\alpha,\beta,\gamma\}=\{\varepsilon,\varepsilon^{\prime},\varepsilon^{\prime\prime}\}$.
The isomorphism classes of simple $\mathbb{C}\mathcal{B}_{n}$-modules are labeled by bipartitions of $n$, i.e., ordered pairs of partitions $(\lambda,\mu)$ such that $|\lambda|+|\mu|=n$. Let $\mathcal{BP}(n)$ denote the set of all bipartitions of $n$. Given $(\lambda,\mu)\in\mathcal{BP}(n)$ with $|\lambda|=a$ and $|\mu|=b$, the corresponding simple $\mathbb{C}\mathcal{B}_{n}$-module $S^{(\lambda,\mu)}$ can be constructed as follows; cf. [8, §5.5.4]. First, the diagonal embedding $GL_{a}(\mathbb{C})\times GL_{b}(\mathbb{C})\subseteq GL_{n}(\mathbb{C})$ induces an embedding $\mathcal{B}_{a}\times\mathcal{B}_{b}\subseteq\mathcal{B}_{n}$. Next, the Specht modules $S^{\lambda}$ and $S^{\mu}$ can be inflated to $\mathcal{B}_{a}$ and $\mathcal{B}_{b}$ via the canonical quotient maps $\mathcal{B}_{a}\twoheadrightarrow\mathfrak{S}_{a}$ and $\mathcal{B}_{b}\twoheadrightarrow\mathfrak{S}_{b}$. Then
@@
If $0<k<n$, then $\binom{n}{k}\geq n$. Now apply (2.2.3) and [10, Theorem 2.4.10]. ∎
@@
By [8, Theorem 5.5.6(b)], the field $\mathbb{Q}$ of rational numbers is a splitting field for $\mathcal{B}_{n}$. We may write $S_{\mathbb{R}}^{(\lambda,\mu)}$ for the simple $\mathbb{R}\mathcal{B}_{n}$-module whose scalar extension to $\mathbb{C}$ is $S^{(\lambda,\mu)}$. Since $\mathbb{Q}$ is a splitting field for $\mathcal{B}_{n}$, it follows that all irreducible characters of $\mathcal{B}_{n}$ are real valued, and hence all simple $\mathbb{C}\mathcal{B}_{n}$-modules are self-dual.
@@ 2.3. Weight space decompositions of Type B simple modules
Our primary reference for this section is [15]. The Young–Jucys–Murphy (YJM) elements $X_{1},\ldots,X_{n}\in\mathbb{C}\mathcal{B}_{n}$ are defined by
@@
Given a partition $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\cdots)$ of $n$, we draw the Young diagram of shape $\lambda$ via the “English” convention (see [18]), as an array of boxes with $\lambda_{i}$ boxes in the $i$-th row, the $(i+1)$-th row drawn below the $i$-th row, and the rows of boxes aligned on the left. A box in the $i$-th row and $j$-th column (from the left) of the diagram is said to have residue $j-i$. A $\lambda$-tableau is a Young diagram of shape $\lambda$ whose boxes have been filled in some order with $n$ different positive integers. A standard $\lambda$-tableau is a $\lambda$-tableau in which the values of the integers increase from top to bottom along columns, and from left to right along rows.
@@
where $\rho_{T}(i)\in\left\{\pm 1\right\}$ is the subscript of the tableau ($T_{+1}$ or $T_{-1}$) in which the integer $i$ is located, and $\operatorname{res}_{T}(i)$ is the residue (or content, in the terminology of [15]) of the box (in either $T_{+1}$ or $T_{-1}$) in which the integer $i$ is located. Then for each $(\lambda,\mu)\in\mathcal{BP}(n)$, one has (cf. [15, Theorem 6.5])
@@ 2.4. Orthogonal forms
The next theorem is a consequence of [15, Theorem 6.12].
@@
    If $i$ and $i+1$ are in the same column of the same tableau in $T$, implying that $r_{i}=-1$, then $s_{i}$ acts on $S_{T}^{(\lambda,\mu)}$ as multiplication by $-1$.
@@
    If $i$ and $i+1$ are in the same row of the same tableau in $T$, implying that $r_{i}=+1$, then $s_{i}$ acts on $S_{T}^{(\lambda,\mu)}$ as multiplication by $+1$.
@@
Given $(\lambda,\mu)\in\mathcal{BP}(n)$, let $R^{(\lambda,\mu)}=(R^{(\lambda,\mu)}_{+1},R^{(\lambda,\mu)}_{-1})$ be the “row major” standard bitableau of shape $(\lambda,\mu)$, in which $R^{(\lambda,\mu)}_{+1}$ (resp. $R^{(\lambda,\mu)}_{-1}$) is filled with the integers $1,\ldots,|\lambda|$ (resp. $|\lambda|+1,\ldots,n$) in row major order (i.e., $R^{(\lambda,\mu)}_{+1}$ contains $1,2,\ldots,\lambda_{1}$ in its first row, then $\lambda_{1}+1,\ldots,\lambda_{2}$ in its second row, etc., and similarly for $R^{(\lambda,\mu)}_{-1}$). Given $T\in\mathbb{T}(\lambda,\mu)$, there exists a unique permutation $\sigma_{T}\in\mathfrak{S}_{n}$ such that $\sigma_{T}\cdot R^{(\lambda,\mu)}=T$. Define the length of $T$ by $\ell(T)=\ell(\sigma_{T})$, the length of $\sigma_{T}$ as an element of the Coxeter group $\mathfrak{S}_{n}$. Then $(-1)^{\ell(T)}=(-1)^{\ell(\sigma_{T})}=\varepsilon^{\prime\prime}(\sigma_{T})$. If $\tau\in\mathfrak{S}_{n}$, then
@@
Say that the action of an adjacent transposition $s_{i}=(i,i+1)$ on a standard bitableau $T\in\mathbb{T}(\lambda,\mu)$ is admissible if $s_{i}\cdot T$ is again a standard bitableau, i.e., if the action corresponds to the first or fourth cases of Lemma 2.4.1. The proof of [15, Lemma 6.2] shows that $T$ can be obtained from the row major standard bitableau $R^{(\lambda,\mu)}$ by applying a sequence of $\ell(T)$ admissible transpositions, i.e., such that at each intermediate step the resulting bitableau is standard. It follows for arbitrary $T,S\in\mathbb{T}(\lambda,\mu)$ that the permutation $\sigma_{T,S}\in\mathfrak{S}_{n}$ that maps $T$ to $S$ can be written as a product of admissible transpositions; let $\ell_{T,S}$ be the minimal number of terms in such a factorization.

### v2

@@ 1.1. The Lie algebra of transpositions
Motivated by questions from the representation theory of the braid group, Marin [Marin:2007] solved the preceding problem in the case $W=\mathfrak{S}_{n}$ of the symmetric group on $n$ letters (i.e., the Weyl group of type $A_{n-1}$) by computing the structure of the Lie algebra $\mathfrak{s}_{n}$ generated by the set of transpositions in $\mathfrak{S}_{n}$. Recall that the isomorphism classes of the simple $\mathbb{C}\mathfrak{S}_{n}$-modules are labeled by integer partitions of $n$. Marin showed that $\mathfrak{s}_{n}$ is a reductive Lie algebra whose semisimple derived subalgebra $\mathfrak{s}_{n}^{\prime}=[\mathfrak{s}_{n},\mathfrak{s}_{n}]$ is a direct sum of special linear Lie algebras, indexed by equivalence classes of partitions whose Young diagrams are non-symmetric and not hooks, and of orthogonal and symplectic Lie algebras, indexed by partitions whose Young diagrams are symmetric and not hooks. A more detailed summary of Marin’s result is given below in Section 4.3.4. The calculation of $\mathfrak{s}_{n}$ enabled Marin to determine the algebraic hull of the image of the braid group in each of the irreducible representations of the Iwahori–Hecke algebra of type A, and in the Iwahori–Hecke algebra itself. Earlier work by Marin [Marin:2003] also showed that the computation of $\mathfrak{s}_{n}$ permits one to describe how tensor products of representations for the Iwahori–Hecke algebra decompose as representations over the braid group; see [BMM:2017, Marin:2013] for related and additional results.
@@ 1.2. The Lie algebra of reflections in types BCD
The proof of Theorem 5.1.1 is by induction on $n$, with the values $n\leq 5$ handled via calculations in GAP [GAP4], and the inductive step for $n\geq 6$ treated in Sections 5.25.10. The overall inductive strategy parallels that employed by Marin for type A, but the statements of results and the arguments required in types B and D can be substantially more involved. In particular, types B and D require an intricate, case-by-case analysis of the implications of the branching rules when one restricts simple modules from $\mathbb{C}\mathcal{B}_{n}$ to $\mathbb{C}\mathcal{B}_{n-1}$ and from $\mathbb{C}\mathcal{D}_{n}$ to $\mathbb{C}\mathcal{D}_{n-1}$.
@@ 1.3. Further questions
The results of this paper open the door to investigating analogues for types BCD of the applications pursued by Marin in type A. For example, one can ask for the algebraic hull of the Artin groups of types B and D inside the corresponding characteristic zero Iwahori–Hecke algebras. The image of these Artin groups inside the finite Iwahori–Hecke algebra (defined over a finite field) has been determined by Esterle [Esterle:2018] in the semisimple case. There may also be applications to algebraic combinatorics: the referee pointed out to us that the representation theory of Lie superalgebras in type A has been used (for example) by Stanley to study unimodality and related properties of partition statistics and other combinatorial sequences [Stanley:1985].
@@
If $W$ is a finite Coxeter group, then there exists a canonical supergroup structure on $W$ (i.e., a multiplicative grading by $\mathbb{Z}/2\mathbb{Z}$) defined by declaring the Coxeter generators to each be of odd superdegree. This grading extends to the group algebra $\mathbb{C}W$, making it an associative superalgebra. One can then consider the induced Lie superalgebra structure on $\mathbb{C}W$, with Lie bracket given by the graded commutator of elements, and ask for the structure of the Lie superalgebra generated by the set of reflections. In [DK:2023] the authors considered this problem in the case $W=\mathfrak{S}_{n}$, and showed that the Lie superalgebra generated by the transpositions is equal to the entire derived Lie superalgebra of $\mathbb{C}\mathfrak{S}_{n}$ plus the span of an odd central element. Generalizing that work to types BCD (and beyond) would be interesting and nontrivial.
@@ 2. Type B classical representation theory
In Sections 2 and 3 we recall some classical results for Weyl groups of types $B_{n}$ and $D_{n}$, many of which can be found in Chapters 1 and 5 of [Geck:2000].
@@ 2.1. Type B Weyl groups
Let $\mathcal{B}_{n}\subseteq GL_{n}(\mathbb{C})$ be the group of signed permutation matrices, i.e., matrices that have precisely one nonzero entry in each row and each column, equal to either +1+1 or 1-1. Let $H_{n}\subseteq\mathcal{B}_{n}$ be the subgroup of permutation matrices in $\mathcal{B}_{n}$ (i.e., matrices in $\mathcal{B}_{n}$ whose nonzero entries are all +1+1), and let $N_{n}\subseteq\mathcal{B}_{n}$ be the subgroup of diagonal matrices in $\mathcal{B}_{n}$. Then $H_{n}$ identifies with the symmetric group $\mathfrak{S}_{n}$, $N_{n}$ identifies with the elementary abelian $2$-group $\mathbb{Z}_{2}^{n}$, and $\mathcal{B}_{n}=N_{n}\rtimes H_{n}$. Henceforward we may make the identification $\mathcal{B}_{n}=N_{n}\rtimes\mathfrak{S}_{n}$.
For $1\leq i\leq n-1$, let $s_{i}\in H_{n}\subseteq\mathcal{B}_{n}$ be the matrix obtained from the identity matrix by interchanging the $i$-th and $(i+1)$-th rows, so $s_{i}$ corresponds to the transposition $(i,i+1)\in\mathfrak{S}_{n}$. For $1\leq i\leq n$, let $t_{i}\in N_{n}\subseteq\mathcal{B}_{n}$ be the diagonal matrix whose $i$-th diagonal entry is equal to 1-1 and whose other diagonal entries are equal to +1+1, so $t_{i}$ corresponds to a multiplicative generator for the $i$-th factor in $\mathbb{Z}_{2}^{n}$. Then for $n\geq 2$, the pair $(\mathcal{B}_{n},\left\{s_{1},\ldots,s_{n-1},t_{n}\right\})$ is a Coxeter system of type $B_{n}$. Each element of $\mathcal{B}_{n}$ is uniquely of the form $t_{1}^{i_{1}}\cdots t_{n}^{i_{n}}\sigma$ with $\sigma\in H_{n}$ and $i_{j}\in\left\{0,1\right\}$ for each $j$. Identifying $H_{n}$ with $\mathfrak{S}_{n}$, the set of reflections in $\mathcal{B}_{n}$ is
@@
The two sets in this union are each conjugacy classes in $\mathcal{B}_{n}$. Following [Okada:1990, §2] or [Mishra:2016, §2], we may label these conjugacy classes as $C_{(21^{n-2},\emptyset)}$ and $C_{(1^{n-1},1)}$, respectively.
@@ 2.2. Simple modules for Type B Weyl groups
The map $\varepsilon$ is the sign character of $\mathcal{B}_{n}$, i.e., the character that evaluates to 1-1 on each of the Coxeter generators of $\mathcal{B}_{n}$. The map $\varepsilon^{\prime\prime}$ factors via the canonical quotient map $\mathcal{B}_{n}=N_{n}\rtimes H_{n}\twoheadrightarrow H_{n}\cong\mathfrak{S}_{n}$ to the usual the sign character on $\mathfrak{S}_{n}$. By abuse of notation, we identify $\varepsilon$, $\varepsilon^{\prime}$, and $\varepsilon^{\prime\prime}$ with their corresponding one-dimensional $\mathbb{C}\mathcal{B}_{n}$-modules. Then $\alpha=\beta\otimes\gamma$ whenever $\{\alpha,\beta,\gamma\}=\{\varepsilon,\varepsilon^{\prime},\varepsilon^{\prime\prime}\}$.
The isomorphism classes of simple $\mathbb{C}\mathcal{B}_{n}$-modules are labeled by bipartitions of $n$, i.e., ordered pairs of partitions $(\lambda,\mu)$ such that $|\lambda|+|\mu|=n$. Let $\mathcal{BP}(n)$ denote the set of all bipartitions of $n$. Given $(\lambda,\mu)\in\mathcal{BP}(n)$ with $|\lambda|=a$ and $|\mu|=b$, the corresponding simple $\mathbb{C}\mathcal{B}_{n}$-module $S^{(\lambda,\mu)}$ can be constructed as follows; cf. [Geck:2000, §5.5.4]. First, the diagonal embedding $GL_{a}(\mathbb{C})\times GL_{b}(\mathbb{C})\subseteq GL_{n}(\mathbb{C})$ induces an embedding $\mathcal{B}_{a}\times\mathcal{B}_{b}\subseteq\mathcal{B}_{n}$. Next, the Specht modules $S^{\lambda}$ and $S^{\mu}$ can be inflated to $\mathcal{B}_{a}$ and $\mathcal{B}_{b}$ via the canonical quotient maps $\mathcal{B}_{a}\twoheadrightarrow\mathfrak{S}_{a}$ and $\mathcal{B}_{b}\twoheadrightarrow\mathfrak{S}_{b}$. Then
@@
If $0<k<n$, then $\binom{n}{k}\geq n$. Now apply (2.2.3) and [James:1981, Theorem 2.4.10]. ∎
@@
By [Geck:2000, Theorem 5.5.6(b)], the field $\mathbb{Q}$ of rational numbers is a splitting field for $\mathcal{B}_{n}$. We may write $S_{\mathbb{R}}^{(\lambda,\mu)}$ for the simple $\mathbb{R}\mathcal{B}_{n}$-module whose scalar extension to $\mathbb{C}$ is $S^{(\lambda,\mu)}$. Since $\mathbb{Q}$ is a splitting field for $\mathcal{B}_{n}$, it follows that all irreducible characters of $\mathcal{B}_{n}$ are real valued, and hence all simple $\mathbb{C}\mathcal{B}_{n}$-modules are self-dual.
@@ 2.3. Weight space decompositions of Type B simple modules
Our primary reference for this section is [Mishra:2016]. The Young–Jucys–Murphy (YJM) elements $X_{1},\ldots,X_{n}\in\mathbb{C}\mathcal{B}_{n}$ are defined by
@@
Given a partition $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\cdots)$ of $n$, we draw the Young diagram of shape $\lambda$ via the “English” convention (see [wiki-Young-tableau]), as an array of boxes with $\lambda_{i}$ boxes in the $i$-th row, the $(i+1)$-th row drawn below the $i$-th row, and the rows of boxes aligned on the left. A box in the $i$-th row and $j$-th column (from the left) of the diagram is said to have residue jij-i. A $\lambda$-tableau is a Young diagram of shape $\lambda$ whose boxes have been filled in some order with $n$ different positive integers. A standard $\lambda$-tableau is a $\lambda$-tableau in which the values of the integers increase from top to bottom along columns, and from left to right along rows.
@@
where $\rho_{T}(i)\in\left\{\pm 1\right\}$ is the subscript of the tableau ($T_{+1}$ or $T_{-1}$) in which the integer $i$ is located, and $\operatorname{res}_{T}(i)$ is the residue (or content, in the terminology of [Mishra:2016]) of the box (in either $T_{+1}$ or $T_{-1}$) in which the integer $i$ is located. Then for each $(\lambda,\mu)\in\mathcal{BP}(n)$, one has (cf. [Mishra:2016, Theorem 6.5])
@@ 2.4. Orthogonal forms
The next theorem is a consequence of [Mishra:2016, Theorem 6.12].
@@
    If $i$ and $i+1$ are in the same column of the same tableau in $T$, implying that $r_{i}=-1$, then $s_{i}$ acts on $S_{T}^{(\lambda,\mu)}$ as multiplication by 1-1.
@@
    If $i$ and $i+1$ are in the same row of the same tableau in $T$, implying that $r_{i}=+1$, then $s_{i}$ acts on $S_{T}^{(\lambda,\mu)}$ as multiplication by +1+1.
@@
[…]
@@
Say that the action of an adjacent transposition $s_{i}=(i,i+1)$ on a standard bitableau $T\in\mathbb{T}(\lambda,\mu)$ is admissible if $s_{i}\cdot T$ is again a standard bitableau, i.e., if the action corresponds to the first or fourth cases of Lemma 2.4.1. The proof of [Mishra:2016, Lemma 6.2] shows that $T$ can be obtained from the row major standard bitableau $R^{(\lambda,\mu)}$ by applying a sequence of $\ell(T)$ admissible transpositions, i.e., such that at each intermediate step the resulting bitableau is standard. It follows for arbitrary $T,S\in\mathbb{T}(\lambda,\mu)$ that the permutation $\sigma_{T,S}\in\mathfrak{S}_{n}$ that maps $T$ to $S$ can be written as a product of admissible transpositions; let $\ell_{T,S}$ be the minimal number of terms in such a factorization.